Using the Rule of 72, at what rate of interest compounded annually will an investment double in five years?

A-13.9%

B-14.4%

C-12.7%

D-15.1%

Using the exact formula, at what rate of interest compounded annually will an investment double in five years?

A-14,17%

B-14.37%

C-14.77%

D-14.87%
Answer #1

Q1
Answer
Option B

Interest rate =72/years=72/5
=14.4%
the interest rate to double the amount in 5 years is 14.4%
---
Q2
ANswer

the exact formula for the interest rate to double the amount is
i=2^(1/n) -1
n=5
i=2^(1/5) -1
i=0.148698355
i=14.87%

the exact interest rate is 14.87%
option D